# Towards Self‑Evolving Benchmarks: Synthesizing Agent Trajectories via Test‑Time Exploration under Validate‑by‑Reproduce Paradigm

### v2

@@ 4.2 Stage 2: Exploration and Trajectory Recording
4.3 Stage 3: Multi-Level Validation of Trajectories
The final stage of our framework is the Trajectory Validator, an autonomous agent that rigorously assesses the quality of evolved tasks. Its method is a step-by-step audit of the solution trace, employing multi-level validation. For each step in the provided solution, the Validator first performs a static analysis, assessing the logical soundness of the thought and verifying that the corresponding code is a faithful implementation of that reasoning. Crucially, this static analysis is complemented by a dynamic execution check to ensure reproducibility. The Validator executes the code from each step itself and compares its own generated output against the observation recorded in the trace. A task only passes this check if the results are semantically equivalent, confirming that the solution is not just logically sound on paper but also practically reproducible. This validation is part of a broader audit that also examines problem integrity, determinism of the answer, and whether the evolved task represents a genuine increase in difficulty. Only tasks that satisfy all criteria, which contains correctness, reproducibility, and increased complexity, are accepted into the final evolved dataset, ensuring a high standard of quality.
Given the inherent randomness of large language model sampling and the loosely coupled nature of our multi-agent system, the generation process can occasionally produce evolved tasks with formatting errors or logical flaws, which consequently fail the rigorous validation stage. To address this, we employ an iterative strategy: multiple evolution attempts are made for tasks in the dataset, but once a task is successfully evolved and validated, it is removed from the candidate pool. This approach allows us to progressively work through the dataset, ensuring that computational resources are focused on tasks yet to be evolved, continuing until a desired proportion of the original dataset has been successfully enhanced.
To further guard against superficially modified but still easy items, we introduce a trajectory-agnostic solver as an auxiliary validator. This solver does not see the generation trajectory; it operates purely under a ReAct (Yao et al., 2023) paradigm and enjoys tool-use parity with the main executor (same multi-modal access, web browsing, and coding). We run the solver under budgeted attempts; if it reliably solves the evolved task within resource limits, the item is flagged as insufficiently challenging and is rejected or sent back for re-evolution. Only tasks that resist this blind, tool-equipped solver proceed, providing an empirical difficulty floor independent of the authored trace. Our pipeline pseudocode can be referred to as Algorithm 1.
@@ 5.1 Experiments Setup
To guarantee the effect of benchmark evolution isolating from model/back-end variance, all evolution stages in Trace use a single back-end, Qwen3-Coder-480B-A35B (Yang et al., 2025), with matched capabilities across agents. Concretely, the Proposer, the Exploration Executor, and the Multi-level Validator are instantiated on the same back-end, receive the same tool-use capability.
For a fair comparison, the only moving part is the LLM back-end; all prompts, decoding configs, tool access, and budgets remain the same. Concretely, we evaluate four LLMs: Kimi K2 (Team et al., 2025), DeepSeek-V3.1 (DeepSeek-AI et al., 2025), Gemini-2.5-Flash (Google DeepMind, 2025), and GPT-5-Mini (OpenAI, 2025). We report per-level and per-round accuracies together with relative deltas from Round 0.
Table 1: Inspiration Emerges: Original vs. Evolved Task by Trace.
Aspect GAIA (Original, Round 0) Evolved by Trace (Round 2 exemplar)
Difficulty Change Low–moderate High
@@ 5.2 Experimental Results and Analysis
5.2.1 Decreased Performance on New Tasks.
@@
AIME-2024
5.2.2 From Seed to Spark.
Candidly, this seed‐to‐spark transformation exceeded our expectations. It demonstrates that Trace can induce benchmark‐level difficulty gains not only by deepening existing reasoning chains, but by rearticulating the task’s background and assessed capability domain through trajectory‐grounded, proposal‐driven autonomous exploration and self‐validation. More cases are shown in Figure 8-15 in our Appendix.
@@ 6 Conclusion
We presented Trace, a self-evolving benchmark framework that couples task evolution with validatable execution trajectories. By separating proposal mining, proposal-guided exploration, and multi-level validation, Trace reliably transforms ceiling-hit benchmarks into harder, trajectory-validatable tasks while preserving the original evaluation interface. On GAIA, successive evolved rounds induce systematic accuracy drops across diverse solvers, indicating genuine benchmark-level difficulty gains rather than isolated item effects.
More broadly, we advocate a paradigm shift from static benchmarks to evolving, dynamics-as-evaluation testbeds. This shift mitigates the community’s dependence on periodically hand-curated “challenging” suites as model capability accelerates under the scaling law: difficulty is refreshed endogenously by the evolution pipeline while remaining reproducible. The recipe is benchmark-agnostic and can be ported to other benchmarks with minimal adaptation, enabling sustainable, auditable, and continually challenging evaluation for agentic systems.
Appendix A The Use of Large Language Models (LLMs)
In this work, we primarily use large language models (LLMs) to aid our writing, including polishing the manuscript by checking grammar errors, and detecting the potential logical flaws.
Appendix B Appendix
In this section, we present the system prompts we set for different agents, as well as examples of tasks before and after evolution.
B.1 Prompts in TRACE framework
B.1.1 The System Prompt of Evolution Proposer Agent
Figure 3 shows our evolution proposal agent’s system prompt. This prompt clearly defines the agent’s tasks and workflow, and provides detailed, categorized hints regarding the agent’s bottlenecks. Guided by this prompt, the agent first analyzes the original task to identify its bottlenecks. Subsequently, it conducts a preliminary exploration of the problem, seeking opportunities to increase its difficulty. Finally, it proposes several feasible proposals. We provided some bottleneck demonstrations for the agent’s reference, as shown in Figure 4.
B.1.2 The System Prompt of Exploration Executor Agent
@@
B.1.3 The Prompt of Trace Validators
@@
B.2 Task Evolution Demonstrations
In this subsection, we present a comparison between pre-evolution and post-evolution problems to demonstrate the effectiveness of our proposed framework. Figure 8-15 show 8 real-world cases of evolution. Specifically, Figures 8, 9, 12, and 13 demonstrate tasks themed around tool invocation and searching, while Figures 10, 11, 14, and 15 present tasks focused on logical reasoning and program writing. For each case, we provide an analysis within the figures, detailing the differences between the two tasks and explaining why the evolved task is more difficult.
@@ B.2 Task Evolution Demonstrations
Appendix Note.
Round Level 1 Level 2 Level 3 Total

### v3

@@ 4.2 Stage 2: Exploration and Trajectory Recording
4.3 Stage 3: Multi-Level Validation of Trajectories
The final stage of our framework is the Trajectory Validator, an autonomous agent that audits evolved tasks through multi-level validation. At the first level, it performs step-by-step replay to ensure reproducibility: for each step in the trajectory, the validator re-executes the tool call and verifies that the resulting output aligns with the recorded observation. At the second level, it conducts a global trace audit to assess the overall logical coherence of the solution. Under the assumption that all intermediate observations are reproducible and the reasoning chain is internally consistent, the task is deemed solvable and well-posed.
Beyond trace validation, the framework evaluates whether the evolved task reflects a genuine increase in difficulty, using the theory-of-mind–inspired bottleneck assessment described earlier. Additional integrity checks are also applied, including answer determinism (to ensure reliable evaluation) and accessibility of referenced resources (e.g., URLs). Only tasks satisfying all criteria including reproducibility, logical validity, increased intrinsic difficulty, and evaluation integrity are accepted into the final dataset.
To further guard against superficially modified but still easy items, we introduce a trajectory-agnostic solver as an auxiliary validator. This solver does not see the generation trajectory and operates purely under a ReAct (Yao et al., 2023) paradigm and enjoys tool-use parity with the main executor (same multi-modal access, web browsing, and coding). We run the solver under budgeted attempts; if it reliably solves the evolved task within resource limits, the item is flagged as insufficiently challenging and is rejected or sent back for re-evolution. Only tasks that resist this blind, tool-equipped solver proceed, providing an empirical difficulty floor independent of the authored trace.
@@ 4.3 Stage 3: Multi-Level Validation of Trajectories
4.4 Test-Time Exploration
Given the inherent stochasticity of large language model sampling and the open-ended nature of trajectory-level evolution, the generation of evolved tasks is not a single-shot procedure, but a structured form of test-time exploration. Within each run, the model engages in multi-turn interactions with real tools and code execution in an open environment, where intermediate states, tool calls, and observations collectively define a candidate solution trajectory from which a new task is constructed.
Crucially, exploration unfolds at two levels. At the intra-run level, the Evolution Executor explores alternative reasoning paths and tool invocations to materialize a trajectory-grounded problem instance. At the inter-run level, multiple such trajectories are generated and subjected to validation. Rather than treating failed attempts as mere sampling noise, we interpret them as natural components of this broader exploration process. Ill-formed, logically inconsistent, or non-reproducible trajectories are filtered through the multi-level validator, which performs step-level replay and global logical auditing. Only trajectories that withstand this scrutiny are admitted into the evolved benchmark.
In this sense, TRACE extends the notion of test-time scaling beyond answer refinement  (Muennighoff et al., 2025): additional computation is allocated not to produce more reliable answers, but to explore and validate harder trajectory-grounded tasks. Thus, test-time scaling becomes a driver of benchmark evolution rather than merely improved problem solving. Our pipeline pseudocode is summarized in Algorithm 1.
@@ 5.1 Experiments Setup
Beyond accuracy, we further track the average token length per task as a proxy for test-time computation and reasoning effort. Concretely, for each model and each round, we fix decoding hyperparameters and compute the mean number of generated tokens over all questions. Intuitively, the more tokens a model needs to produce before committing to an answer, the longer it must stay in the loop to perform additional reasoning steps, tool calls, and self-corrections.
@@
To guarantee the effect of benchmark evolution isolating from model/back-end variance, all evolution stages in Trace use a single back-end, Qwen3-Coder-480B-A35B (Yang et al., 2025), with matched capabilities across agents. Concretely, the Proposer, the Exploration Executor, and the Multi-level Validator are instantiated on the same back-end, receive the same tool-use capability. The auxiliary validator model is Qwen3-235B-A22B-Instruct (Yang et al., 2025).
For a fair comparison, the only moving part is the LLM back-end; all prompts, decoding configs, tool access, and budgets remain the same. Concretely, we evaluate four LLMs: Kimi-K2 (Team et al., 2025), DeepSeek-V3.1 (DeepSeek-AI et al., 2025), Gemini-2.5-Flash (Google DeepMind, 2025), and GPT-5-Mini (OpenAI, 2025). We report per-level and per-round accuracies together with relative deltas from Round 0.
@@ 5.2 Experimental Results and Analysis
5.2.1 Longer Reasoning, Lower Pass@1 on evolved tasks
@@
As shown in Table 2, average answer length increases for GPT-5-mini and KIMI-K2 substantially as the benchmark evolves, while Pass@1 concurrently degrades. This joint trend indicates that TRACE is not merely introducing noise, but is making tasks genuinely harder in a way that forces models into longer, more demanding reasoning trajectories.
@@
Table 2: Token length results on the GAIA-TRACE benchmark.
Metric GPT-5-Mini KIMI-K2
Evo. $\leftarrow$ Orig. $\Delta$ Evo. $\leftarrow$ Orig. $\Delta$
Avg. Length (Round 1) $\mathbf{4898.2}\leftarrow 2864.5$ $\mathbf{(+2033.7)}$ $\mathbf{6609.0}\leftarrow 3389.6$ $\mathbf{(+3219.4)}$
Avg. Length (Round 2) $\mathbf{6275.7}\leftarrow 2864.5$ $\mathbf{(+3411.2)}$ $\mathbf{8454.7}\leftarrow 3389.6$ $\mathbf{(+5065.1)}$
AIME-2024
@@
Table 4: Inspiration Emerges: Original vs. Evolved Task by Trace.
Aspect GAIA (Original, Round 0) Evolved by Trace (Round 2 exemplar)
Difficulty Change Low–moderate High
5.2.2 From Seed to Spark.
Candidly, this seed‐to‐spark transformation exceeded our expectations. It demonstrates that TRACE can induce benchmark‐level difficulty gains not only by deepening existing reasoning chains, but by rearticulating the task’s background and assessed capability domain through trajectory‐grounded, proposal‐driven autonomous exploration and self‐validation. More cases are shown in Figure 9-16 in our Appendix.
@@ 6 Conclusion
We introduce Trace, a framework that simulates human expert workflows to autonomously evolve benchmarks, replacing labor-intensive manual curation. Through bottleneck analysis that guides exploration in real-world environments, Trace transforms saturated tasks into structurally harder, trajectory-validated challenges. Empirically, it raises difficulty barriers on benchmarks such as GAIA, exposing model limitations often hidden by static evaluation. Trace promotes a shift toward sustainable, self-evolving benchmarks that remain rigorous, reproducible, and aligned with rapidly advancing agent capabilities.
Acknowledgement
We gratefully acknowledge the support from Shanghai Artificial Intelligence Laboratory.
Appendix
In this section, we present extended experimental results and visualizations. We also provide the prompts we set for different agents, as well as examples of tasks before and after evolution.
Extended Experimental Results and Visualizations
Table 5 shows how TRACE evolution reshapes the capability mix of GAIA. Compared to the original dataset, the evolved benchmark slightly down-weights pure web-browsing and multi-modality questions, while increasing the share of coding and diverse filetype reading tasks, which typically require more structured reasoning and robust tool use. The higher “Integrative Capabilities” mass further reflects the emergence of more complex, mixed-capability items that do not fall cleanly into a single category, indicating that TRACE places greater emphasis on multi-faceted agentic behavior rather than simple lookup-style queries.
Table 5: Comparison of capability distribution ratios before and after TRACE evolution on GAIA.
Capability Category GAIA Original (%) Evolve Round 1 (%) Capability Totals (%)
Web browsing 43.9 42.2 27.1
Coding 19.1 20.5 24.7
Multi-modality 17.1 4.8 8.2
Diverse filetype reading 16.0 20.5 28.2
Integrative Capabilities 4.0 12.0 11.8
Figure 3: Capability-wise profiles explaining ranking shifts on TRACE-evolved GAIA. For each model, radar plots show accuracy across capability categories, together with the distribution of tasks in each category (outer pink region), on Round 1 (top) and Round 2 (bottom). TRACE evolution re-weights the benchmark distribution, while simultaneously demanding longer, tool-heavy trajectories, which in turn exposes different capability “fingerprints” across models and accounts for changes in their relative ranking.
Round Level 1 Level 2 Level 3 Total
Prompts in TRACE framework
The System Prompt of Evolution Proposer Agent
Figure  4 shows our evolution proposal agent’s system prompt. This prompt clearly defines the agent’s tasks and workflow, and provides detailed, categorized hints regarding the agent’s bottlenecks. Guided by this prompt, the agent first analyzes the original task to identify its bottlenecks. Subsequently, it conducts a preliminary exploration of the problem, seeking opportunities to increase its difficulty. Finally, it proposes several feasible proposals. We provided some bottleneck demonstrations for the agent’s reference, as shown in Figure 5.
The System Prompt of Exploration Executor Agent
@@ Acknowledgement
The Prompt of Trace Validators
@@
Task Evolution Demonstrations
In this subsection, we present a comparison between pre-evolution and post-evolution problems to demonstrate the effectiveness of our proposed framework. Figure 9-16 show 8 real-world cases of evolution. Specifically, Figures 9, 10, 13, and 14 demonstrate tasks themed around tool invocation and searching, while Figures 11, 12, 15, and 16 present tasks focused on logical reasoning and program writing. For each case, we provide an analysis within the figures, detailing the differences between the two tasks and explaining why the evolved task is more difficult.